# Near Coincidences and Nilpotent Division Fields

### v1

@@ 1.2. Acknowledgements
We would like to thank David Zureick-Brown and Álvaro Lozano-Robledo for helpful conversations during the writing of this paper, and we would like to thank Pedro Lemos for helpful communications about the results in [20].
@@ 2.1. Elliptic Curves
$\ \bar{\rho}_{E,n}\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$
with the property that $\operatorname{Im}\bar{\rho}_{E,n}\simeq\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q}).$ We remark here that because the isomorphism $E[n]\simeq\mathbb{Z}/n\mathbb{Z}\oplus\mathbb{Z}/n\mathbb{Z}$ is non-canonical, $\operatorname{Im}\bar{\rho}_{E,n}$ is really only defined up to conjugation. This will not be a major point throughout the paper, but is worth pointing out.
@@
[34, Lemma 2] Let $p$ be a prime and let $G$ be a subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. If $p$ divides $|G|$ then, either $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})\subseteq G,$ or $G$ is contained inside a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. If $p$ does not divide $|G|$, let $H$ be the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$; then
@@
[28, Section IV] Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 5$ be a prime. If $\bar{\rho}_{E,p}$ is surjective, then $\rho_{E,p^{\infty}}$ is also surjective.
@@
  • $G$ contains an element of determinant $-1$ and trace 00 that fixes a point of order $n$ inside of $(\mathbb{Z}/n\mathbb{Z})^{2}.$
@@
[36, Proposition 2.2] Let $n\geq 2$ be an integer and let $E/\mathbb{Q}$ be an elliptic curve. Then $\operatorname{Im}\bar{\rho}_{E,n}$ is an admissible subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$.
@@ 2.1.1. Elliptic curves with complex multiplication
Elliptic curves come in two distinct types depending on their endomorphism rings. Given an elliptic curve $E$, defined over a field of characteristic 0, the endomorphism ring of $E$ over $\overline{\mathbb{Q}}$ is either isomorphic to $\mathbb{Z}$ or an order of an quadratic imaginary field, usually denoted by $\mathcal{O}$. When the endomorphism ring is larger than $\mathbb{Z}$ we say that $E$ has complex multiplication by $\mathcal{O}$. Throughout this section we follow the work done in [26, Section 12]. Many of the results we use were first proven in [22]. A reader looking for an introduction to elliptic curves with complex multiplication should see [30, Chapter II].
@@
$\mathcal{C}_{\mathcal{O}}(\widehat{\mathbb{Z}})=\lim_{\leftarrow}\mathcal{C}_{\mathcal{O}}(N)\subseteq\operatorname{GL}_{2}(\widehat{\mathbb{Z}}).$
The group $\mathcal{C}_{\mathcal{O}}(\widehat{\mathbb{Z}})$ is a closed subgroup of $\operatorname{GL}_{2}(\widehat{\mathbb{Z}})$ that is isomorphic to $\mathcal{C}_{\mathcal{O}}$ under the isomorphism
@@
and let $\mathcal{N}_{\mathcal{O}}(\widehat{\mathbb{Z}})\subseteq\operatorname{GL}_{2}(\widehat{\mathbb{Z}})$ and $\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})\subseteq\operatorname{GL}_{2}(\widehat{\mathbb{Z}}_{p})$ be the usual inverse limits of the $\mathcal{N}_{\mathcal{O}}(N)$.
@@
Frequently the group $\mathcal{N}_{\mathcal{O}}$ is called the normalizer of $\mathcal{C}_{\mathcal{O}}.$ It turns out that the group $\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$ is the normalizer of $\mathcal{C}_{\mathcal{O}}(\mathbb{Z}_{p})$ in $\operatorname{GL}_{2}(\mathbb{Z}_{p})$ (See [22, Proposition 5.6(2)]), but $\mathcal{N}_{\mathcal{O}}$ is not the normalizer of $\mathcal{C}_{\mathcal{O}}$ in $\operatorname{GL}_{2}(\widehat{\mathbb{Z}}).$ See [26, Remark 12.1.2].
Before moving on we make a few more observations about these groups that will be useful later on. First we note that by construction each of the groups $\mathcal{C}_{\mathcal{O}}(N)$ is an abelian group since $\mathcal{O}/N\mathcal{O}$ is abelian. In contrast, the groups $\mathcal{N}_{\mathcal{O}}$ are not abelian unless $p=2$. In order to compute the center of $\mathcal{N}_{\mathcal{O}}(N)$, it would a simple matter of determining which matrices in $\mathcal{C}_{\mathcal{O}}(N)$ commute with $M=\left(\begin{smallmatrix}-1&0\\ \phi&1\end{smallmatrix}\right).$ Let
@@
Computing the entry in the first row, second column of $MA$ and $AM$ we see that $A$ commutes with $M$ if and only if $b=-b$. Thus, if $p\neq 2$, then $A\in Z(\mathcal{N}_{\mathcal{O}}(N))$ if and only if $A=\begin{pmatrix}a&0\\ 0&a\end{pmatrix}=aI\in Z(\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})).$ From this we get the following lemma.
@@
The next lemma we need is about the sizes of these groups in $p$-adic towers.
Lemma 2.7.
Let $\mathcal{O}$ be an order of a quadratic imaginary field $K$, let $p$ be a prime and let $k\geq 2$ be a positive integer. Then
$|\mathcal{C}_{\mathcal{O}}(p^{k})|=p^{2(k-1)}|\mathcal{C}_{\mathcal{O}}(p)|,\hbox{and consequently }|\mathcal{N}_{\mathcal{O}}(p^{k})|=p^{2(k-1)}|\mathcal{N}_{\mathcal{O}}(p)|.$
Proof.
Consider the map $\pi\colon\mathcal{C}_{\mathcal{O}}(p^{k})\to\mathcal{C}_{\mathcal{O}}(p)$ given by component-wise reduction. The kernel of this map is exactly the set of
$\begin{pmatrix}a+b\phi&b\\ \delta b&a\end{pmatrix}\in\mathcal{C}_{\mathcal{O}}(N)$
congruent to the identity $\bmod\;p$. Looking at the second column, this occurs exactly when $a\equiv 1\bmod p$ and $b\equiv 0\bmod p$. So, the kernel of this map has size $(p^{k-1})^{2}$ and the result follows from the first isomorphism theorem. ∎
The last few theorems we need will help us pin down the exact image of the Galois representations attached to elliptic curves with complex multiplication. We will state the following theorems for elliptic curves with complex multiplication defined over $\mathbb{Q}$, but we note that both [26, Section 12] and [22] handle the case where $E$ is defined over a number field.
@@
[26, Proposition 12.1.4] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathcal{O}$, let $p$ be a prime, and let
@@
Lemma 2.7 and Proposition 2.8 will be exactly what we need in order to understand how the division fields change as we go up the $p$-adic tower. This will be useful in Section 5.
@@
[22, Theorem 1.2(4)] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathcal{O}\neq\mathbb{Z}[\zeta_{3}].$ If $p$ does not divide $2\operatorname{disc}(\mathcal{O})$, then there is a choice of basis such that $\operatorname{Im}\rho_{E,p^{\infty}}=\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$.
@@ 2.2. Modular Curves
The nature of this correspondence depends on whether $-I\in G$ or not, but if $G$ were a nilpotent group that did not contain $-I$, then adding $-I$ would preserve nilpotency. For this reason, we can assume that $-I\in G$. Then the curve $X_{G}$ always comes with a natural morphism
@@
such that an elliptic curve $E/\mathbb{Q}$ with $j$-invariant $j_{E}\not\in\{0,1728\}$, has $\operatorname{Im}\bar{\rho}_{E,n}$ conjugate to a subgroup of $G$ if and only if $j_{E}=\pi_{G}(P)$ for some $P\in X_{G}(\mathbb{Q})$. The interested reader should see [26, Subsection 2.3] to see what happens when $-I\not\in G$.
@@
3. Near Coincedences
Before starting the proof in earnest, we will start by taking care of the case when $E$ has complex multiplication. The base case will break down into a number of subcases. We will have to handle the cases when $p=2$ and $3$ separately, but even further in the case when $p\geq 5,$ we will have to break this into cases depending on $\operatorname{Im}\bar{\rho}_{E,p}$ according to Proposition 2.1. If $E/\mathbb{Q}$ is an elliptic curve with complex multiplication and $p\geq 5$ then $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$ cannot hold since by Lemma 2.7 and Proposition 2.8 the extension $\mathbb{Q}(E[p^{2}])/\mathbb{Q}(E[p])$ must be degree $p^{2}$. If $E/\mathbb{Q}$ is an elliptic curve with complex multiplication and $p=2$ or 3, then we can use the fact that all of the possible images of $\rho_{E,p^{\infty}}$ are listed in [26, Tables 18–22]. With this information, we can completely answer the question of near coincidences of division fields for elliptic curves with complex multiplication.
@@ 3. Near Coincedences
Let $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$ with surjective determinant and suppose that $m\mid n$. We say that $G$ represents a near coincedence of level $(n,m)$ if
@@
where $\pi:\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})\to\operatorname{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$ is the standard componentwise reduction map.
@@
We will start the case when $n=1$ of the theorem by dealing with primes $p\geq 5$ and break the argument into cases depending on $\operatorname{Im}\bar{\rho}_{E,p}$ based on Proposition 2.1.
In the first case, if $\operatorname{Im}\bar{\rho}_{E,p}=\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$, then we know that $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$ since $\operatorname{Im}\bar{\rho}_{E,p^{2}}=\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ by Proposition 2.2. In the second case, when $\operatorname{Im}\bar{\rho}_{E,p}$ is contained inside of a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$, we know that $E$ has a cyclic $p$-isogeny. Given this, we can use the classification of cyclic $p$-isogenies together with the data gathered in [26] to rule out the possibility that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}}).$ In the next case, when $\operatorname{Im}\bar{\rho}_{E,p}$ is contained inside of the normalizer of a split Cartan, we will make use of a small result as well as the results of [3].
@@
Let $E/\mathbb{Q}$ be an elliptic curve and $p\geq 3$ a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is conjugate to a subgroup of a normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ and $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}}).$ Then, $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ is conjugate to a subgroup of a normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$.
@@
To start the proof, we let $\pi\colon\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})\to\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ be the standard componentwise reduction map and let $G=\operatorname{Im}\bar{\rho}_{E,p}.$ A priori, we know that $\operatorname{Im}\bar{\rho}_{E,p^{2}}\subseteq\pi^{-1}(G)$. Further, since the size of $G$ is a divisor of $2(p-1)^{2}$ and $|\pi^{-1}(G)|=p^{4}|G|,$ we know that $\pi^{-1}(G)$ has a $p$-complement isomorphic to $G$. In fact, if we let $(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}=\langle\alpha\rangle$, then this $p$-complement is a subgroup of
$\left\langle\begin{pmatrix}\alpha^{p}&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 0&\alpha^{p}\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\rangle.$
We also note that since $p\geq 3$, it must be that this $p$-complement must be non-trivial.
Next, if we suppose that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}}),$ we can use the fact that $\mathbb{Q}(E[p])\cap\mathbb{Q}(\zeta_{p^{2}})=\mathbb{Q}(\zeta_{p})$ to decompose $\operatorname{Gal}(\mathbb{Q}(E[p^{2}])/\mathbb{Q})$. In particular,
$\operatorname{Gal}(\mathbb{Q}(E[p^{2}])/\mathbb{Q})\simeq\{(\sigma,\tau)\in\operatorname{Gal}(\mathbb{Q}(E[p])/\mathbb{Q})\times\operatorname{Gal}(\mathbb{Q}(\zeta_{p^{2}})/\mathbb{Q})\colon\sigma\mid_{\mathbb{Q}(\zeta_{p})}=\tau\mid_{\mathbb{Q}(\zeta_{p})}\}.$
Let $\tau_{0}\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p^{2}})/\mathbb{Q})$ be an element of order $p$ that fixes $\mathbb{Q}(\zeta_{p})$. Such an element exists since $\operatorname{Gal}(\mathbb{Q}(\zeta_{p^{2}})/\mathbb{Q})\simeq(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}\simeq\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/(p-1)\mathbb{Z}.$ Then, there is an element of $\operatorname{Gal}(\mathbb{Q}(E[p^{2}])/\mathbb{Q})$ of order exactly $p$ corresponding to $(Id,\tau_{0})$, call this element $\eta$. Clearly, by construction, $\eta$ is in the center of $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ and thus must commute with all the elements of $\operatorname{Gal}(\mathbb{Q}(E[p^{2})/\mathbb{Q}).$ Further, $\operatorname{Im}\bar{\rho}_{E,p^{2}}(\eta)$ must commute with the $p$-complement above. Notice in this case, that if $\mathbb{Q}(E[p])/\mathbb{Q}$ is an abelian extension, then so is $\mathbb{Q}(E[p^{2}]/\mathbb{Q})$, but according to [14] this cannot happen for $p\geq 3.$ Thus, it must be that not only is this $p$-complement non-trivial, it is also non-abelian. That is, it must contain $A=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and an element that is not of the form $aI$ of $aA$.
To summarize, we know that the $p$-complement is non-abelian and we know that $\bar{\rho}_{E,p^{2}}(\eta)$ must commute with everything in the $p$-complement. In particular, $\bar{\rho}_{E,p^{2}}(\eta)$ commutes with $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and either matrix of the form $\begin{pmatrix}a&0\\ 0&b\end{pmatrix}$ or $\begin{pmatrix}0&a\\ b&0\end{pmatrix}$. The only way that this can happen is if $\bar{\rho}_{E,p^{2}}(\eta)$ is of the form $aI$ for some $a\in(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}$. This forces $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ to be a subgroup of a normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z}).$
So, in order to determine if one can have $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$ when $p\geq 5$ and $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a split Cartan, we can use the the classification of elliptic curves that at have their mod $p^{2}$ images contained in the normalizer of a split Cartan subgroup from [3] to see that this doesn’t happen.
The last case that has to be dealt with is the case when $\operatorname{Im}\bar{\rho}_{E,p}$ is contained inside of the normalizer of a non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ for $p\geq 5$. One important fact in this case is that any such elliptic curve has to have potentially supersingular reduction at $p$ from either [18, Appendix B] or [36, Proposition 1.13].
@@
[29] Let $E/\mathbb{Q}$ be an elliptic curve with potential good reduction at $p\geq 5$ and discriminant $\Delta$. Then, $E$ acquires good reduction over $\mathbb{Q}(\sqrt[12]{\Delta})$ at all primes over $p$.

### v3

@@ 1.2. Acknowledgements
We would like to thank David Zureick-Brown and Álvaro Lozano-Robledo for insightful conversations during the writing of this paper, and we would like to thank Pedro Lemos for helpful communications about the results in [25]. We would also like to thank the anonymous referees whose careful reading has helped improve the clarity and accuracy of the paper.
@@ 2.1. Elliptic Curves
[…]
with the property that $\operatorname{Im}\bar{\rho}_{E,n}\simeq\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q}).$ We remark here that because the isomorphism $E[n]\simeq\mathbb{Z}/n\mathbb{Z}\oplus\mathbb{Z}/n\mathbb{Z}$ is non-canonical, $\operatorname{Im}\bar{\rho}_{E,n}$ is really only defined up to conjugation.
@@
[42, Lemma 2] Let $p$ be a prime and let $G$ be a subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. If $p$ divides $|G|$ then, either $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})\subseteq G,$ or $G$ is contained inside a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. If $p$ does not divide $|G|$, let $H$ be the image of $G$ in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$; then
@@
[35, Section IV] Let $E/\mathbb{Q}$ be an elliptic curve and let $p\geq 5$ be a prime. If $\bar{\rho}_{E,p}$ is surjective, then $\rho_{E,p^{\infty}}$ is also surjective.
@@
  • $G$ contains an element of determinant 1-1 and trace 0 that fixes a point of order $n$ inside of $(\mathbb{Z}/n\mathbb{Z})^{2}.$
@@
[45, Proposition 2.2] Let $n\geq 2$ be an integer and let $E/\mathbb{Q}$ be an elliptic curve. Then, $\operatorname{Im}\bar{\rho}_{E,n}$ is an admissible subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$.
@@ 2.1.1. Elliptic curves with complex multiplication
Elliptic curves come in two distinct types depending on their endomorphism rings. Given an elliptic curve $E$, defined over a field of characteristic 0, the endomorphism ring of $E$ over $\overline{\mathbb{Q}}$ is either isomorphic to $\mathbb{Z}$ or an order of an imaginary quadratic field, usually denoted by $\mathcal{O}$. When the endomorphism ring is larger than $\mathbb{Z}$ we say that $E$ has complex multiplication by $\mathcal{O}$. Throughout this section we follow the work done in [33, Section 12]. Many of the results we use were first proven in [28]. A reader looking for an introduction to elliptic curves with complex multiplication should see [38, Chapter II].
@@
[…]
@@
[…]
@@
[…]
Before moving on we make a few more observations about these groups that will be useful later on. First we note that by construction each of the groups $\mathcal{C}_{\mathcal{O}}(N)$ is an abelian group since $\mathcal{O}/N\mathcal{O}$ is abelian. In contrast, the groups $\mathcal{N}_{\mathcal{O}}$ are not abelian unless $p=2$. In order to compute the center of $\mathcal{N}_{\mathcal{O}}(N)$, it would be a simple matter of determining which matrices in $\mathcal{C}_{\mathcal{O}}(N)$ commute with $M=\left(\begin{smallmatrix}-1&0\\ \phi&1\end{smallmatrix}\right).$ Let
@@
Computing the entry in the first row, second column of $MA$ and $AM$ we see that $A$ commutes with $M$ if and only if b=bb=-b. Thus, if $p\neq 2$, then $A\in Z(\mathcal{N}_{\mathcal{O}}(N))$ if and only if $A=\begin{pmatrix}a&0\\ 0&a\end{pmatrix}=aI\in Z(\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z})).$ From this we get the following lemma.
@@
The last few theorems we need will help us pin down the exact image of the Galois representations attached to elliptic curves with complex multiplication. We will state the following theorems for elliptic curves with complex multiplication defined over $\mathbb{Q}$, but we note that both [33, Section 12] and [28] handle the case where $E$ is defined over a number field.
@@
[33, Proposition 12.1.4] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathcal{O}$, let $p$ be a prime, and let
@@
Proposition 2.7 will be exactly what we need in order to understand how the division fields change as we go up the $p$-adic tower. This will be useful in Section 5.
@@
[28, Theorem 1.2(4)] Let $E/\mathbb{Q}$ be an elliptic curve with complex multiplication by $\mathcal{O}\neq\mathbb{Z}[\zeta_{3}].$ If $p$ does not divide $2\operatorname{disc}(\mathcal{O})$, then there is a choice of basis such that $\operatorname{Im}\rho_{E,p^{\infty}}=\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$.
@@ 2.2. Modular Curves
The nature of this correspondence depends on whether $I\in G$ or not, but if $G$ were a nilpotent group that did not contain I-I, then adding I-I would preserve nilpotency. For this reason, we can assume that $I\in G$. Then the curve $X_{G}$ always comes with a natural morphism
@@
such that an elliptic curve $E/\mathbb{Q}$ with $j$-invariant $j_{E}\not\in\{0,1728\}$, has $\operatorname{Im}\bar{\rho}_{E,n}$ conjugate to a subgroup of $G$ if and only if $j_{E}=\pi_{G}(P)$ for some $P\in X_{G}(\mathbb{Q})$. The interested reader should see [33, Subsection 2.3] to see more details when $I\not\in G$.
@@
3. Near coincidences
@@
Let $G\subseteq\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$ with surjective determinant and suppose that $m\mid n$. We say that $G$ represents a near coincidence of level $(n,m)$ if
@@
where $\pi\colon\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})\to\operatorname{GL}_{2}(\mathbb{Z}/m\mathbb{Z})$ is the standard componentwise reduction map.
@@
We will start the case when $n=1$ of the theorem by dealing with primes $p\geq 5$ and break the argument into cases depending on $\operatorname{Im}\bar{\rho}_{E,p}$ based on Proposition 2.1. First we handle the case that $\bar{\rho}_{E,p}$ is surjective.
@@
Assume that $E/\mathbb{Q}$ is an elliptic curve, $p\geq 5$ is prime, and $\operatorname{Im}\bar{\rho}_{E,p}=\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
@@
By Proposition 2.2, we have that $\operatorname{Im}\rho_{E,p^{\infty}}=\operatorname{GL}_{2}(\mathbb{Z}_{p})$ and hence $\operatorname{Im}\bar{\rho}_{E,p^{2}}=\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$. Hence
$\mathbb{Q}(E[p^{2}]):\mathbb{Q}]=|\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})|=p^{5}(p-1)^{2}(p+1),$
while
$[\mathbb{Q}(E[p],\zeta_{p^{2}})]$ $\leq[\mathbb{Q}(E[p],\zeta_{p^{2}}):\mathbb{Q}(E[p])][\mathbb{Q}(E[p]):\mathbb{Q}]\leq p(p-1)\cdot|\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})|$
$=p^{2}(p-1)^{3}(p+1)<p^{5}(p-1)^{2}(p+1).$
This proves the claim. ∎
Next we handle the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is a proper subgroup whose order is a multiple of $p$.
@@
Assume that $E/\mathbb{Q}$ is an elliptic curve, $p\geq 5$ is prime, and $\operatorname{Im}\bar{\rho}_{E,p}$ is a proper subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ whose order is a multiple of $p$. Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
Proof.
In this case, the image of $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in a Borel subgroup by Proposition 2.1 and hence $E$ has a cyclic $p$-isogeny defined over $\mathbb{Q}$. By the classification of Mazur, we have that $p\in\{5,7,11,17,19,43,67,163\}$. Let $H=\mathrm{im}\,\rho_{E,p^{\infty}}$.
If $E$ has complex multiplication, choose a basis so that $H\subseteq\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$. By Proposition 2.7, $H$ is the full preimage in $\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$ of $\operatorname{Im}\bar{\rho}_{E,p}$. Since $\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$ contains all matrices of the form $\begin{bmatrix}1+kp&0\\ 0&1-kp\end{bmatrix}$ with $k\in\mathbb{Z}/p\mathbb{Z}$ and this implies that $\mathbb{Q}(E[p^{2}])/\mathbb{Q}(E[p],\zeta_{p^{2}})$ is a non-trivial extension.
If $E$ does not have complex multiplication, then Theorem 1.1.6 of [33] implies that $X_{H}$ is isomorphic to $\mathbb{P}^{1}$ or a positive rank elliptic curve, or that $X_{H}$ is listed in Table 1 of [33] as the other cases are ruled out by the assumption that $p\mid|\operatorname{Im}\bar{\rho}_{E,p}|$. If we assume that $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$, then $p^{2}\nmid|\mathrm{im}\,\bar{\rho}_{E,p^{2}}|$ and this implies that the index of $H$ in $\operatorname{GL}_{2}(\mathbb{Z}_{p})$ is a multiple of $p^{3}$. Since $p\geq 5$, this implies that the index is $\geq 125$ and this implies that $X_{H}$ has genus $\geq 2$. There are groups $H$ contained in a Borel subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ listed in Table 1 of [33] for $p=11$, $17$ and $37$, but in no case is the index of $H$ a multiple of $p^{3}$. This concludes the proof. ∎
Next we handle the case that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a split Cartan subgroup. Here we divide the argument into the CM case and the non-CM case.
Proposition 3.5.
Let $E/\mathbb{Q}$ be an elliptic curve and $p\geq 5$ a prime such that $\operatorname{Im}\bar{\rho}_{E,p}$ is contained in the normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. Then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$.
Proof.
We first will show that $E$ has complex multiplication. If $E$ does not have complex multiplication, then the results of [3] rule out $p=11$ and $p\geq 17$, and the results of [1] rule out $p=13$. Hence $p\in\{5,7\}$. If $H=\operatorname{Im}\bar{\rho}_{E,p^{2}}$ and $\mathbb{Q}(E[p^{2}])=\mathbb{Q}(E[p],\zeta_{p^{2}})$, then $H$ is an index $p^{3}$ subgroup of the full preimage in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ of $\operatorname{Im}\bar{\rho}_{E,p}$ and hence $|H|$ divides $\frac{|\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})|}{p^{4}(p+1)/2}$, since $\frac{p(p+1)}{2}$ is the index of the normalizer of the split Cartan in $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$. In addition, by Proposition 2.4, $H$ must also be an admissible group. Finally, $H$ must be non-abelian, since by [16] $\mathbb{Q}(E[n])/\mathbb{Q}$ is non-abelian if $n>8$.
For these two primes $p$, we use Magma to search for admissible subgroups $H$ of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$ so that (i) the order of $H$ divides $\frac{|\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})|}{p^{4}(p+1)/2}$, and (ii) $H$ is non-abelian. For $p=5$ there are two conjugacy classes of such subgroups and for $p=7$ there are four. In both cases, all such subgroups are contained in the normalizer of a split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p^{2}\mathbb{Z})$. (The code for these calculations can be found in the file prop35.m at [8].) The rational points on the modular curve corresponding to the normalizer of the split Cartan subgroup mod $25$ are determined by Momose and Shimura in [29], and the mod $49$ case is handled by Momose in [30]. In both cases, all such elliptic curves have CM. It follows that $E$ has complex multiplication.
Now we show that if $E$ has complex multiplication, then $\mathbb{Q}(E[p^{2}])\neq\mathbb{Q}(E[p],\zeta_{p^{2}})$. The here is identical to that in the proof of Proposition 3.4. Choose a basis so that $\operatorname{Im}\bar{\rho}_{E,p^{2}}\subseteq\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$. By Proposition 2.7, $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ is the full preimage in $\mathcal{N}_{\mathcal{O}}(\mathbb{Z}_{p})$ of $\operatorname{Im}\bar{\rho}_{E,p}$. Therefore, $\operatorname{Im}\bar{\rho}_{E,p^{2}}$ contains all matrices of the form $\begin{bmatrix}1+kp&0\\ 0&1-kp\end{bmatrix}$ with $k\in\mathbb{Z}/p\mathbb{Z}$ and this implies that $\mathbb{Q}(E[p^{2}])/\mathbb{Q}(E[p],\zeta_{p^{2}})$ is a non-trivial extension. ∎
The last case that has to be dealt with is the case when $\operatorname{Im}\bar{\rho}_{E,p}$ is contained inside of the normalizer of a non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ for $p\geq 5$. One important fact in this case is that any such elliptic curve has to have potentially supersingular reduction at $p$ from either [22, Appendix B] or [45, Proposition 1.13].
Proposition 3.6.
@@
[36, p. 312] Let $E/\mathbb{Q}$ be an elliptic curve with potential good reduction at $p\geq 5$ and discriminant $\Delta$. Then, $E$ acquires good reduction over $\mathbb{Q}(\sqrt[12]{\Delta})$ at all primes over $p$.